# Scaling Laws for Energy Efficiency of Local LLMs

### v3

@@ I Introduction
Edge AI is expanding across robotics, medical devices, drones, personal computing, and industrial systems, and these deployments are dominated by CPU-only hardware, particularly low-power ARM architectures. However, academic and industrial research overwhelmingly benchmarks models on GPUs, and often in data-center settings [green_ai2020, henderson2020, mlperf_v40]. This leaves a critical open question: what are the computational requirements and scaling laws of large language–model and vision–language–model inference when they are deployed as local assistants on CPUs?
While recent work has begun to develop inference- and test-time scaling laws for large language and multimodal models [wu2024_inference_scaling, chen2025_inference_scaling, bansal2024_testtime_scaling, fu2024_vlm_compute_optimal], these studies characterize performance in terms of abstract compute (e.g., FLOPs) or GPU-accelerated inference. Existing research on edge and on-device AI primarily focuses on throughput, latency, or quantization quality, and very few studies establish empirical scaling laws describing how CPU-only inference cost evolves with input complexity (e.g., token length or image resolution) [edge_ai_survey2023, previous_edge_llm2023, cpu_inference2023]. In parallel, the literature on energy-aware and “green AI” benchmarks has begun to formalize measurement protocols and metrics for efficiency [green_ai_measurement2023, ipw2025], yet often still assumes GPU-accelerated or heterogeneous platforms.
@@
Energy-aware evaluation has gained prominence in the context of ”green AI” and carbon accounting. Lacoste et al. formalized methodologies for quantifying the carbon emissions of machine learning workloads [green_ai_measurement2023], and Saad-Falcon et al. introduced the Intelligence-per-Watt (IPW) benchmark to compare local AI systems under efficiency constraints [ipw2025]. Our work is complementary: rather than defining a new efficiency benchmark, we investigate how CPU-only multimodal inference behaves as a function of input complexity and compression, and we provide per-prompt and per-run energy measurements on representative consumer and embedded hardware that are commonly used for local LLM inference.
@@
On the systems side, llama.cpp and the GGUF format have become de facto standards for efficient on-device deployment of LLMs and VLMs on commodity CPUs [llama_cpp2024, gguf_format2024]. Our experiments build directly on this ecosystem, using llama.cpp both as a runtime and as a controllable preprocessing pipeline for VLMs. For CPU/RAM profiling, prior work has advocated integrated metrics based on area under resource-usage curves [auc_computation2022], which capture both intensity and duration of utilization; we adopt this AUC perspective as our main system-level cost metric. Energy measurements on edge systems commonly rely on USB-based or board-level power meters [usb_power_measurement2023], and we follow this line with an inline USB-C power meter and per-prompt energy accounting.
@@ II-C Resource and energy measurement
CPU and RAM usage are sampled at 5 Hz, and we compute an Area Under Curve (AUC) above a stable idle baseline via the trapezoidal rule, following prior work on resource-usage profiling [auc_computation2022]. We sample CPU and RAM at 5 Hz because higher frequencies introduced substantial high-frequency noise, while lower frequencies lacked the temporal resolution needed to reliably detect inference windows. The AUC we report is not the ML-classification metric but the time-integrated resource usage: the x-axis is time (seconds) and the y-axis is CPU or RAM utilization above a stable idle baseline. Thus, the AUC measures total computational effort over the duration of the prompt. AUC integrates total system effort across time, abstracting away transient bursts and normalizing for multi-core differences. Energy is measured using an inline USB-C power meter (1 Hz sampling), consistent with prior methodologies for edge AI power measurement [usb_power_measurement2023]. All metrics are computed per-prompt, and inference windows are automatically detected from the gradient of the CPU signal. While higher sampling frequencies would reduce integration error, our method follows common practice in edge-AI energy studies.
@@ IV Discussion
Energy and sustainability also matter at deployment time. Our measurements show that LLM energy drops by 50–62% and VLM energy by 6–37% under compression, in line with the broader push toward green AI and carbon-aware evaluation [green_ai_measurement2023, ipw2025]. At scale, even modest per-call savings accumulate, especially when combined with realistic electricity prices such as those reported by Eurostat for European households [eurostat_electricity_prices_2024]. Taken together, the results on MacBook Pro M2 and Raspberry Pi 5 indicate that, with appropriate compression and preprocessing, local LLMs can operate efficiently across a wide range of CPU-only devices, from high-end laptops to low-cost edge boards.

### v4

@@ I Introduction
Edge AI is expanding across robotics, medical devices, drones, personal computing, and industrial systems, and these deployments are dominated by CPU-only hardware, particularly low-power ARM architectures. However, academic and industrial research overwhelmingly benchmarks models on GPUs, and often in data-center settings [green_ai2020, henderson2020]. This leaves a critical open question: what are the computational requirements and scaling laws of large language–model and vision–language–model inference when they are deployed as local assistants on CPUs?
While recent work has begun to develop inference- and test-time scaling laws for large language and multimodal models [wu2024_inference_scaling, chen2025_inference_scaling, bansal2024_testtime_scaling, fu2024_vlm_compute_optimal], these studies characterize performance in terms of abstract compute (e.g., FLOPs) or GPU-accelerated inference. Existing research on edge and on-device AI primarily focuses on throughput, latency, or quantization quality, and very few studies establish empirical scaling laws describing how CPU-only inference cost evolves with input complexity (e.g., token length or image resolution) [edge_ai_survey2023, previous_edge_llm2023, cpu_inference2023]. In parallel, the literature on energy-aware and “green AI” benchmarks has begun to formalize measurement protocols and metrics for efficiency [green_ai_measurement2019, ipw2025], yet often still assumes GPU-accelerated or heterogeneous platforms.
@@
Energy-aware evaluation has gained prominence in the context of ”green AI” and carbon accounting. Lacoste et al. formalized methodologies for quantifying the carbon emissions of machine learning workloads [green_ai_measurement2019], and Saad-Falcon et al. introduced the Intelligence-per-Watt (IPW) benchmark to compare local AI systems under efficiency constraints [ipw2025]. Our work is complementary: rather than defining a new efficiency benchmark, we investigate how CPU-only multimodal inference behaves as a function of input complexity and compression, and we provide per-prompt and per-run energy measurements on representative consumer and embedded hardware that are commonly used for local LLM inference.
@@
On the systems side, llama.cpp and the GGUF format have become de facto standards for efficient on-device deployment of LLMs and VLMs on commodity CPUs [llama_cpp2024, gguf_format2024]. Our experiments build directly on this ecosystem, using llama.cpp both as a runtime and as a controllable preprocessing pipeline for VLMs. For CPU/RAM profiling, prior work has advocated integrated metrics based on area under resource-usage curves [sklearn_auc], which capture both intensity and duration of utilization; we adopt this AUC perspective as our main system-level cost metric. Energy measurements on edge systems commonly rely on USB-based or board-level power meters [wang2024_power_estimation], and we follow this line with an inline USB-C power meter and per-prompt energy accounting.
@@ II-C Resource and energy measurement
CPU and RAM usage are sampled at 5 Hz, and we compute an Area Under Curve (AUC) above a stable idle baseline via the trapezoidal rule, following prior work on resource-usage profiling. We sample CPU and RAM at 5 Hz because higher frequencies introduced substantial high-frequency noise, while lower frequencies lacked the temporal resolution needed to reliably detect inference windows. The AUC we report is not the ML-classification metric but the time-integrated resource usage: the x-axis is time (seconds) and the y-axis is CPU or RAM utilization above a stable idle baseline. Thus, the AUC measures total computational effort over the duration of the prompt. AUC integrates total system effort across time, abstracting away transient bursts and normalizing for multi-core differences. Energy is measured using an inline USB-C power meter (1 Hz sampling), consistent with prior methodologies for edge AI power measurement. All metrics are computed per-prompt, and inference windows are automatically detected from the gradient of the CPU signal. While higher sampling frequencies would reduce integration error, our method follows common practice in edge-AI energy studies.
@@ IV Discussion
Energy and sustainability also matter at deployment time. Our measurements show that LLM energy drops by 50–62% and VLM energy by 6–37% under compression, in line with the broader push toward green AI and carbon-aware evaluation [green_ai_measurement2019, ipw2025]. At scale, even modest per-call savings accumulate, especially when combined with realistic electricity prices such as those reported by Eurostat for European households [eurostat_electricity_prices_2024]. Taken together, the results on MacBook Pro M2 and Raspberry Pi 5 indicate that, with appropriate compression and preprocessing, local LLMs can operate efficiently across a wide range of CPU-only devices, from high-end laptops to low-cost edge boards.